OpenStudy (anonymous):

What is the value of x? Drawing will be in the comments shortly. (Note: F = 93.5 degrees.)

OpenStudy (anonymous):

 Drawing is not to scale.

OpenStudy (anonymous):

I'll just assign the value y to angle DEF 93.5 + x + y = 180 x + y = 86.5 since line DF and line EF are equidistant, the triangle is isosceles, meaning that x = y x + y = 86.5 x + x = 86.5 I reckon you can do it from here.

OpenStudy (anonymous):

I got 43.25. Rounded it would make 43.3. Am I correct?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Awesome, thank you! :)